Hi
4/7 of a rectangular tank with a base area of 2000 cm^2 was filled with water. Another 6.9 litres of water were needed to fill up the tank until it was 9/10 full.
What was the height of the tank
What is the final volume of water in the tank

Answer by mananth(16946)

4/7 of a rectangular tank with a base area of 2000 cm^2 was filled with water. Another 6.9 litres of water were needed to fill up the tank until it was 9/10 full.
What was the height of the tank
What is the final volume of water in the tank
Let height of tank be h
Volume of tank = 2000h (area * h)
Volume of water in tank = 4/7 * 2000h
6.9 litres was added = 6900 cm^3
After adding 6.9 litres 9/10 of tank is full
Relation
4/7 * 2000h +6900 = 9/10 * 2000h
9/10 * 2000h-4/7 * 2000h= 6900
Rearrange
2000h (9/10- 4/7)= 6900
2000h*23/70 = 6900
Rearrange
(70*6900)/(23*2000)=h
=10.5 cm
Final volume of water
= (9/10)* 2000*10.5 =18900 cm^3
Convert to litres if required